OpenStudy (anonymous):

Suppose that L = 2, R = 40, E(t) = 100e^-10t, and I(0) = 0. Find the maximum current in the circuit for t >= 0.

OpenStudy (anonymous):

is it parallel or series ?

OpenStudy (anonymous):

I don't think that information is relevant here.

OpenStudy (anonymous):

Ldi/dt+Ri=E(t) 2di/dt+40i=100e-10t di/dt+20i=50e-10t i(t)=ce-20t+5e-10t i(0)=0=c+5->c=-5 i(t)=-5e-20t+5e-10t di/dt=0-> 100e-20t=50e-10t e-10t=1/2-> e-20t=1/4 imax=-5/4+5/2=5/4

OpenStudy (anonymous):

Wait. I(t) = ce^-20t + 5e^-10t. Where does that come about?

OpenStudy (anonymous):

y'+ay=b(t) y=ce-at+e-at(int e^at*b(t))

OpenStudy (anonymous):

Suppose a is a constant.

OpenStudy (anonymous):

Wouldn't the particular solution Ip = Ke^-10t

OpenStudy (anonymous):

yes for particular form you can choose k and then find it out.